OpenStudy (iwanttogotostanford):

(x^2-4x+3)-(3x^2-4x-3) is equivalent to: 2x^2-6 2x^2-8x 2x^2-8x-6 -2x^2+6 -2x^2-8x

OpenStudy (iwanttogotostanford):

@Vocaloid

Vocaloid (vocaloid):

start by distributing the negative sign

OpenStudy (iwanttogotostanford):

how do i do that

Vocaloid (vocaloid):

hint: -(A-B-C) = -A+B+C

OpenStudy (iwanttogotostanford):

oh ok gotcha

Vocaloid (vocaloid):

-(3x^2-4x-3) = ?

OpenStudy (iwanttogotostanford):

(3x^2+-4x+-3)?

Vocaloid (vocaloid):

almost, check your signs

OpenStudy (iwanttogotostanford):

oh, -3x^2+-4x+-3?

Vocaloid (vocaloid):

getting closer

Vocaloid (vocaloid):

-(-4x) = ? -(-3) = ?

OpenStudy (iwanttogotostanford):

4x and 3!

Vocaloid (vocaloid):

right, so we have - 3x^2 + 4x + 3 on the right side after we distribute now we put the whole problem back together

OpenStudy (iwanttogotostanford):

yes

Vocaloid (vocaloid):

x^2-4x+3- 3x^2 + 4x + 3 know where to go from here?

OpenStudy (iwanttogotostanford):

just combine right?

Vocaloid (vocaloid):

yes, so what does that give us?

OpenStudy (iwanttogotostanford):

x^2+14-3x^2? Im really bad at combining like terms for some reason..

OpenStudy (iwanttogotostanford):

i most likely did it wrong

Vocaloid (vocaloid):

not quite x^2 - 3x^2 = ?

OpenStudy (iwanttogotostanford):

3

OpenStudy (iwanttogotostanford):

so it would be x^+11?

Vocaloid (vocaloid):

uh, not quite hint: 1 - 3 = -2 x - 3x = -2x x^2 - 3x^2 = ?

OpenStudy (iwanttogotostanford):

just -2?

Vocaloid (vocaloid):

don't forget the x^2

OpenStudy (iwanttogotostanford):

-2^2???

OpenStudy (iwanttogotostanford):

im so bad at this one ...

Vocaloid (vocaloid):

nope, when you combine like terms, you keep the variable x^2 - 3x^2 = -2x^2

OpenStudy (iwanttogotostanford):

plus 6 right?

Vocaloid (vocaloid):

right

Vocaloid (vocaloid):

think of it this way: 1 - 3 = -2 1 dollar - 3 dollars = -2 dollars the "dollar" stays in the answer, just like x^2 stays in the answer

Vocaloid (vocaloid):

so our final answer is -2x^2 + 6

OpenStudy (iwanttogotostanford):

ok, im gonna need to practice that one a bit. thanks for all of you help! is it ok if i tag you in a couple of questions tomorrow?

Vocaloid (vocaloid):

sure, I'm online all day tomorrow

OpenStudy (iwanttogotostanford):

ok- thanks i will tag you !